# Estimation of Food Intake Quantity Using Inertial Signals from Smartwatches

### v1

@@ I INTRODUCTION
Recent research focuses on solutions employing commercial devices to reduce intrusiveness and hardware complexity. Particularly, 3D accelerometer and gyroscopes, commonly embedded in commercial smartwatches, have demonstrated the interesting potential in dietary monitoring applications. Kyritsis et al. developed a deep learning framework for modeling meal micromovements using commercial smartwatch data [10], successfully classifying fundamental eating gestures into distinct movement patterns. Their subsequent work extended these capabilities to unrestricted environments [11], demonstrating robust performance in both bite detection and meal boundary identification across free-living conditions. While these advances establish the effectiveness of smartwatch IMU data for detecting eating episodes, their potential for quantitative measures like bite weight estimation remains unexplored.
@@ ACKNOWLEDGMENT
This work was partially supported by REBECCA (European Union’s Horizon 2020 research and innovation programme under grant agreement No 965231) and BETTER4U (European Union’s Horizon Europe research and innovation programme under grant agreement No 101080117).

### v2

@@ I INTRODUCTION
Recent research focuses on solutions employing commercial devices to reduce intrusiveness and hardware complexity. Particularly, IMU sensors, commonly embedded in commercial smartwatches, have demonstrated the interesting potential in dietary monitoring applications. Kyritsis et al. developed a deep learning framework for modeling meal micromovements using commercial smartwatch data [10], successfully classifying fundamental eating gestures into distinct movement patterns. Subsequent work extended these capabilities to free-living environments, demonstrating robust bite detection, meal boundary identification, and micro-behavior analysis in both healthy individuals and those with Parkinson’s Disease [11, 12]. While these advances establish the effectiveness of smartwatch IMU data for detecting eating episodes, their potential for quantitative measures like bite weight estimation remains unexplored.
@@ ACKNOWLEDGMENT
This work was partially supported by REBECCA (European Union’s Horizon 2020 research and innovation programme under grant agreement No 965231).